Tom buys $20,000 worth of stock, and sells it eight months later for $28,000. If he is required to pay 15% capital gains tax on his profits, how much tax will he pay?
Answer by orca(409)

Note that the capital gain tax = profit*rate
The profit = 28,000 - 20,000 = 8,000
The tax to be paid = 8,000*15% = 1,200